Operations Report – February 9th



Crew 274 Operations Report 2/9/2023

SOL: 11

Name of person filing report: Alexis Joseph Lojek

Non-nominal systems: Toilet. Robotic Observatory.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice.


Spirit rover used: Yes

Hours: 208.8

Beginning charge: 100%

Ending charge: 53%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.9

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Curiosity rover used: No

Hours: 220.7

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: 255.4

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: Rovers Spirit and Opportunity were used for the morning EVA and plugged in when they returned, as it was a cloudless day and before noon. Curiosity and Perseverance were not utilized.

Summary of Hab operations: Toilet tank was emptied this evening and habitat is in a nominal condition.

WATER USE: 42 gallons as of 6:50pm

Water (static tank): 236 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged after EVA today and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 12:39 – 1:04 pm.

WATER USE: 14.79 gallons

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: Yes; 86 grams of tomatoes. Total: 86 grams.

Summary of ScienceDome operations: GHO Tyler Hines’ experiment concluded today. The Microgreens were a “partial success” according to Hines, with cress being the most successful to germinate. The LDEF seeds – one that was previously noted as germinating on Sol 6 is still germinating, and will be preserved for future study. Th rest of the seeds that did not germinate will be discarded.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The MDRS Robotic Observatory is inoperable until further notice due to the mechanical issue described in Sol 9’s Operations report.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site